Problem 3. Write the summation with the first two terms and the last term outside of the summation. Use the sequences from 1(a) and 1(d) to complete this problem. 10 j-2 m-+2 3(b) Let m 20. dk
Answer #1

Hey,

Note: Brother In case of any queries brother just comment in box I would be very happy to assist all your queries.

Below is the answer

a)

Since ai is GP with a=8 and r=2

So, the nth term of sequuence is a*r^n-1

So, the summation with first and last 2 outside are

So, summation=

=

b)

Since kth term is first k+1 positive odds

So, d0=1

d1=1+3=4

d_(m+1)=1+3+5+.... (m+2 terms)

((m+2)/2)*(2+2(m+1))=(m+2)^2

d_(m+2)=1+3+5+....(m+2 terms)

((m+3)/2)*(2+2*(m+2))=(m+3)^2

So, the summation is


Kindly revert for any queries

Thanks.